A plate of thickness t behaves as a half wave plate for a light of wavelength λ. ignoring variation in refractive indices with λ, if a light of 2λ is used then plate will behave as? (i know answer is quarter wave plate. But i don't know why... So please explain me)
1
Expert's answer
2020-06-08T10:29:27-0400

The thickness of a half-wave plate for wavelength λ is given by:


"t=\frac{λ}{2(\mu_O-\mu_E)}=\frac{2λ}{4(\mu_O-\mu_E)}=\frac{λ'}{4(\mu_O-\mu_E)}"

Thus, for wavelength λ' = 2λ, it will behave as a quarter-wave plate.